ow many grams of solute are present in 495 mL of 0.700 M KBr? NumberI multiplied the concentration of solute x volume of solution molar mass and got = 41241.123 grams; this does not seem right.
Answer #1

Molarity = (W / MW) (1000 / V in mL)

Molarity = 0.700 M   

MW = 119 g/mol

V = 495 mL

W = ?

0.700 = (W / 119) (1000 / 495)

W / 119 = 0.3465

W = 41.23 g

answer = 41.23 g